OpenStudy (anonymous):

How do I determine the domain from a graph model, in which it gives reasonable results?

OpenStudy (primeralph):

This is the region in which the x values lie. Post an example and I'll show you how.

OpenStudy (primeralph):

You need the domain of which graph?

OpenStudy (anonymous):

The Yellow (parabola) and dark blue (catenary)

OpenStudy (primeralph):

Now you need to know the model of what you're making. For example, say you were plotting distance against time, the left half will be useless because it corresponds to negative time, which is not reasonable. So what models are these for?

OpenStudy (anonymous):

I'm researching the conflict between the two mathematical models for a hanging necklace chain. Catenary- y=a/2[e^bx+e^-bx] and parabola- y=y=ax^2+bx+c..

OpenStudy (anonymous):

And Discuss the domain within which both models give reasonable results..

OpenStudy (primeralph):

Well, that depends on your goals. Really can't help you unless I fully know what you're trying to accomplish.

OpenStudy (anonymous):

I traced the curve of a chain, (light necklace) onto a number place. Then found the equation of the parabol that closely approximates the curve I drew. Then I derived an equation for the catenary. Then put all, the original, parabola, and catnerary on the same graph to compare which fits the original curve. This is the catenary. Now I am to discuss its validity and the domain which both models give reasonable results. I am unsure of how to find the domain..

OpenStudy (primeralph):